OpenStudy (anonymous):

The perimeter of a rectangle is 42 cm. the length is 3 cm greater than the width. Find the width.

OpenStudy (anonymous):

Set up an equation. 42= w+w+(3+w)+(3+w) 42= 4w+6 42-6= 4w+6-6 36=4w 36/4=4w/4 9 = w So w=9; The width is 9 cm.